Suppose the block of the previous problem s pushed down the incline with a force of 4N.how far will the block move in the first two seconds after starting from rest?the mass of the block is 4kg.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Net acceleration would be = gsinthetha + 4/4 = > gsin(thetha) + 1 m/s^2.
In first two seconds,
s = 0 + 0.5*(gsin(thetha) + 1)*2*2 = > 2gsinthetha + 2 meter.
thetha is the angle whatevr the incline plane has in your previous problem.
As your previous problem cannot be visualised here.
Hence distance it would travel is:
 2gsinthetha + 2 meter.